OpenStudy (anonymous):

|2x-1|+|x+3|=5.......... find all real numbers

OpenStudy (anonymous):

I2x-1I+Ix+3I=5 I2x-1I=5-Ix+3I

OpenStudy (anonymous):

now square both sides

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

one moment im just working through it

OpenStudy (anonymous):

sorry im just learning i dont know where to take the square

OpenStudy (anonymous):

I2x-1I+Ix+3I=5 I2x-1I=5-Ix+3I Squaring both sides gives, (2x-1)^2= (5-Ix+3I)^2 4x^2 -4x+1=25-10Ix+3I+x^2+6x+9 3x^2-10x-33=-10Ix+3I

OpenStudy (anonymous):

Now, can u?

OpenStudy (anonymous):

yes thanks